A certain population is known to be growing at a rate given by the logistic equation dx/dt=x(b-ax),where a and b are positive constants .
Show that the maximum rate of growth will occur when the population is equal to half the equilibrium size, that is, when the population is (b/2a).
1
Expert's answer
2020-09-27T18:12:10-0400

The logistic equation is given by "\frac{dx}{dt}=x(b-ax)" where "\frac{dx}{dt}" is the rate of growth .


Let "y=\frac{dx}{dt}" ..so we have to show that y is maximum when the population "(x) = \frac{b}{2a}" .


Now

"y= x(b-ax)"


Now differentiation both sides w.r.t "x," we get

"\frac{dy}{dx}=x(-a)+(b-ax)*1 \space" (using product rule of differentiation.)


"or, \frac{dy}{dx}=b-2ax\\"


for maxima /minima we must have "\frac{dy}{dx}=0" ,( to get the critical points)


"\therefore b-2ax = 0\\or, x= \frac{b}{2a}\\ Now \space \frac{d^2y}{dx^2}= -2a \\i.e \space \frac{d^2y}{dx^2}\space is\space negative \space ( because \space a \space is \space positive \space constant )\\ \therefore y \space is \space maximum \space at \space x=\frac{b}{2a}"


Hence the maximum rate of growth will occur when the population is equal to half the equilibrium size, that is, when the population is (b/2a).


Hence PROVED.